# Learning to Discover Generalized Facial Expressions

### v1

@@ 2.3.2 Fine-tuning of Main Classification Head
Table 2: Category discovery accuracy (ACC) on FerPlus.
Sur.+Fea.+Dis.+Hap. Sur.+Fea.+Dis.+Ang. Sur.+Hap.+Sad.+Neu.
Method All Old New All Old New All Old New
k-means [13] 20.8 22.1 20.2 22.5 23.0 22.3 25.1 25.6 24.0
GCD [7][CVPR’22] 40.8 62.3 27.4 40.4 52.2 32.2 37.2 50.4 29.7
GPC [8][ICCV’23] 45.2 63.7 34.4 45.7 64.3 40.0 52.2 58.5 46.8
InfoSieve [14][NeurIPS’23] 48.2 74.1 37.7 47.5 72.5 44.1 64.1 69.5 53.7
SimGCD [10][ICCV’23] 45.8 65.4 38.3 46.4 70.2 42.2 63.7 67.2 55.0
Ours 51.8 84.9 38.8 48.6 81.7 43.5 68.6 74.3 56.4
Table 3: Category discovery accuracy (ACC) on large-scale dataset AffectNet.
Sur.+Fea.+Dis.+Hap. Sur.+Fea.+Dis.+Ang. Sur.+Hap.+Sad.+Neu.
Method All Old New All Old New All Old New
k-means [13] 14.4 14.6 14.3 15.0 20.3 13.2 14.2 15.2 13.8
GCD[7][CVPR’22] 35.0 47.5 28.1 32.2 42.0 28.9 28.0 38.2 24.6
GPC [8][ICCV’23] 36.3 46.5 33.4 38.7 50.5 32.1 33.6 44.1 30.0
InfoSieve [14][NeurIPS’23] 47.1 54.2 35.7 49.8 68.4 37.2 46.2 57.3 31.4
SimGCD [10][ICCV’23] 45.6 53.4 35.1 40.7 55.1 34.5 38.1 48.1 29.7
Ours 57.7 62.9 40.9 54.2 74.8 39.4 55.9 69.4 30.8
3 Experiments
3.1 Experimental Setup
Dataset. We conduct experiments on three popular FER datasets, which have been partitioned in the format suitable for GCD. RAF-DB [16]contains eleven compound expressions and seven basic expressions. For GCD, we treat four out of the seven basic expressions as known categories, with three different selection methods. Notably, in the stage of discovering compound expressions, we use the basic expressions as known categories. FERPlus [17] is extended from FER2013 [18]. We used eight categories and selected four as known categories to discover the remaining new categories. AffectNet [19] is a large-scale FER dataset. We used eight categories and selected four as known categories to discover the remaining new categories. A summary of dataset statistics is shown in the appendix B.1. Please refer to the appendix for a statistical summary of the dataset and the reasons for the specific category partitioning.
Implementation details. For a fair comparison, we ran our method and the comparison methods five times on each dataset, and ultimately reported the set of results where the "All" accuracy ranked third. We employ a ViT-B/16 backbone network [21] pre-trained with DINO [22]. Training was performed using an initial learning rate of 0.1, which was decayed with a cosine annealed schedule [23] . The max training epoch $T$ is set to 200 and batch size of 128 for training. We follow SimGCD [10] to set the balancing factor $\lambda$ to 0.35, and the temperature values $\tau_{c}$ and $\tau_{u}$ to 0.1 and 0.07, respectively. We initially set $\tau_{t}$ to 0.07 and $\tau_{s}$ to 0.1 for the classification objective. Then, a cosine schedule is employed to gradually reduce $\tau_{t}$ to 0.04 over the first 30 epochs. $\lambda_{a}$, $\lambda_{b}$, and $\lambda_{c}$ are assigned values of 0.2, 0.3, and 0.2, respectively. $T_{warmup}$ is set to 50 epochs. The coefficients $\alpha$ and $\beta$ are set to 0.2 and 0.1, respectively. The All experiments are conducted using an NVIDIA GeForce RTX 3090 GPU.
Figure 4: Comparison with open-set FER [5] in the open-set scenario.
Table 4: Ablation studies of different components of our method on the RAF-DB basic expression classes, showing both implicit and explicit debiasing effects.
Sur.+Fea.+Dis.+Hap. Sur.+Fea.+Dis.+Ang.
Method All Old New All Old New
SimGCD (baseline) 43.3 47.9 40.2 49.6 56.9 48.5
w/ $\mathcal{L}_{\text{ad}}$ (implicit debias only) 50.4+7.1 63.2+15.3 40.5+0.3 50.5+0.9 60.5+3.6 48.6+0.1
w/o $\mathcal{L}_{\text{ad}}$ (explicit debias only) 49.8+6.5 61.1+13.2 40.7+0.5 51.5+1.9 63.2+6.3 50.3+1.8
w/o $\mathcal{L}_{\text{cluster}}$ (implicit + $\mathcal{L}_{\text{bal}}$) 52.3+9.0 65.3+17.4 40.4+0.2 50.7+1.1 63.4+6.5 47.5-1.0
w/o $\mathcal{L}_{\text{bal}}$ (implicit + $\mathcal{L}_{\text{cluster}}$) 51.4+8.1 64.2+16.3 40.7+0.5 50.7+1.1 61.5+4.6 48.9+0.4
Ours (all components) 53.3+10.0 71.3+23.4 41.0+0.8 52.4+2.8 66.3+9.4 50.2+1.7
Figure 5: Attention visualization of different heads (numbered as h1 to h3) on RAF-DB. The top 10% attended patches are shown in red. Our method pays more attention to the cheeks, eyes and mouth corners details. Our model learns more discriminative features than other methods.
3.2 Comparison with the State-of-the-Art methods
Comparision with FER method. Given that the previous FER methods were not tailored for the GCD task, we opted for Ada-CAM [15], a semi-supervised approach, as a comparative baseline. We subjected Ada-CAM to our GCD evaluation metrics, and the results, as shown in Tab. 1, underscore the superiority of our method in the category discovery task. Furthermore, in light of the recent emergence of the open-set scenario for FER, introduced by [5] et al., we adapted the evaluation metrics for FER-GCD to align with the Open-set paradigm. Specifically, we utilized AUROC (higher is better) and FPR@TPR95 (lower is better) as our evaluation metrics. As shown in Fig. 4, FER-GCD achieved higher AUROC metrics, representing the greater capability for new category detection. In addition, the low FPR@TPR95 also reflects the confidence level of FER-GCD in making decisions, further reflecting the effectiveness of the debiasing framework.
Comparision with GCD methods. The effectiveness of FER-GCD debiasing is most directly reflected in its ability to maintain high accuracy for old facial expression categories while also achieving excellent results in discovering new facial expression categories. In Tab. 1, we present a comparison with the SOTA GCD method on the basic classes of the RAF-DB dataset. In terms of old category accuracy, we achieved an average improvement of 16.2% compared to SimGCD [10]. And the ability to discover new expressions has increased steadily. In Tab. 6, we report the comparison on the challenging RAF-DB-Compound dataset that discovers composite expressions based on basic expressions, we achieved an improvement of 20.8% accuracy for the old categories and 4.9% for the new categories, compared to SimGCD. Moreover, Tab. 2 shows the results on the dataset FerPlus [17]. Compared to SimGCD, we achieved significant improvements, whereas InfoSeive [14] is able to achieve an even higher accuracy for new categories when Surprise, Fear, Disgust, and Anger were known categories. Furthermore, Tab. 3 shows performance comparison on the large-scale dataset AffectNet [19]. On this challenging dataset, we both achieve the best new category, old category accuracy. In summary, our approach effectively eliminates the bias towards new categories while maintaining high accuracy for old categories, and steadily improves the recognition accuracy for new categories. An important basis for the outstanding performance of our method in identifying old categories and discovering new categories is that it forms more discriminative feature boundaries, as illustrated in Fig. 8.
3.3 Ablation Study
In this section, we analyze the importance of the components we introduce mainly on the RAF-DB basic classes.
Effect of estimating the maximum bias (implicit debias). We first examine the role of estimating maximum bias, employing an auxiliary classification head for adversarial estimation to ensure robust performance of the feature extractor under maximum bias. Tab.  4 presents the ablation results. We strive to minimize interference from new category introduction on old categories while improving accuracy for new categories.
Enhancing the effect of clustering (explicit debias). In FER-GCD, we introduce the clustering algorithm after the warmup epoch $T_{\text{warmup}}$, assuming that the model has preliminary clustering capabilities by then. This allows us to further constrain the feature space through our clustering algorithm. Tab.  4 presents the ablation results.
Focus on minority categories (explicit debias). A smoother dynamic focus mechanism allows us to focus well on a small number of categories and further improves overall accuracy. Tab.  4 demonstrates the validity of this technique.
4 Visualisation
As shown in Fig. 5, we adopt the method from GCD [7] to conduct a visual analysis of the attention maps generated by the DINO-ViT model. Specifically, we visualize the attention Heads 1 through 3, which provides valuable insights into the crucial aspects of the model’s recognition process, particularly how it captures the subtle nuances in facial expressions. Besides, the t-SNE visualization results shown in Fig. 8 demonstrate a more discrete feature representation of FER-GCD, compared to the more blurred feature boundaries of GCD and SimGCD.
5 Conclusion
In this paper, we introduce Generalized Category Discovery (GCD) for facial expression recognition (FER), with the aim of reducing manual annotation costs and improving the model’s ability to detect new compound expressions, thereby enhancing future sentiment analysis. During training, interference between old and new classes can bias the model, especially in fine-grained and imbalanced FER scenarios. To address this, we first theoretically defined the bias and proposed a new framework called FER-GCD, which debiases through implicit and explicit steps. Finally, our method is able to maintain good performance in recognizing old expressions while enhancing the ability to discover new expressions. Our work aims to advance facial expression recognition for open-world tasks.
Appendix A Theory
A.1 Notation

### v2

@@ 4.3 Overall Loss Function:
DIG-FACE 53.3 71.3 41.0 52.4 66.3 50.2 69.1 77.8 41.9
Table 3: Category discovery accuracy (ACC) on FerPlus and AffectNet datasets.
FerPlus AffectNet
Sur.+Fea.+Dis.+Hap. Sur.+Fea.+Dis.+Ang. Sur.+Hap.+Sad.+Neu. Sur.+Fea.+Dis.+Hap. Sur.+Fea.+Dis.+Ang. Sur.+Hap.+Sad.+Neu.
Method All Old New All Old New All Old New All Old New All Old New All Old New
k-means [27] 20.8 22.1 20.2 22.5 23.0 22.3 25.1 25.6 24.0 14.4 14.6 14.3 15.0 20.3 13.2 14.2 15.2 13.8
GCD [35] 40.8 62.3 27.4 40.4 52.2 32.2 37.2 50.4 29.7 35.0 47.5 28.1 32.2 42.0 28.9 28.0 38.2 24.6
GPC [48] 45.2 63.7 34.4 45.7 64.3 40.0 52.2 58.5 46.8 36.3 46.5 33.4 38.7 50.5 32.1 33.6 44.1 30.0
InfoSieve [30] 48.2 74.1 37.7 47.5 72.5 44.1 64.1 69.5 53.7 47.1 54.2 35.7 49.8 68.4 37.2 46.2 57.3 31.4
CMS-GCD [5] 48.9 77.6 36.5 46.5 78.5 40.1 62.7 75.2 47.8 - - - - - - - - -
SimGCD [39] 45.8 65.4 38.3 46.4 70.2 42.2 63.7 67.2 55.0 45.6 53.4 35.1 40.7 55.1 34.5 38.1 48.1 29.7
DIG-FACE 51.8 84.9 38.8 48.6 81.7 43.5 68.6 74.3 56.4 57.7 62.9 40.9 54.2 74.8 39.4 55.9 69.4 30.8
5 Experiments
5.1 Experimental Setup
Dataset. We conduct experiments on three popular FER datasets, which have been partitioned in the format suitable for G-FACE. RAF-DB  [21] contains eleven compound expressions and seven basic expressions. For G-FACE, we treat four out of the seven basic expressions as known categories, with three different selection methods. Notably, in the stage of discovering compound expressions, we use the basic expressions as known categories. FERPlus [2] is extended from FER2013 [11]. We used eight categories and selected four as known categories to discover the remaining new categories. AffectNet [28] is a large-scale FER dataset. We used eight categories and selected four as known categories to discover the remaining new categories. A summary of dataset statistics is shown in the appendix. Please refer to the appendix for a statistical summary of the dataset and the reasons for the specific category partitioning.
Implementation details. For a fair comparison, we conducted the comparison methods five times on each dataset, and ultimately reported the set of results where the ”All” accuracy ranked third. We employed a ViT-B/16 backbone (feature exactor $\psi$) network [7] pre-trained with DINO [4] (decision head). Training was performed using an initial learning rate of 0.1, which was decayed with a cosine annealed schedule [25]. The max training epoch $T$ is set to 200 and batch size of 128 for training. We followed SimGCD [39] to set the balancing factor $\lambda$ to 0.35, and the temperature values $\tau_{c}$ and $\tau_{u}$ to 0.1 and 0.07, respectively. We initially set $\tau_{t}$ to 0.07 and $\tau_{s}$ to 0.1 for the classification objective. Then, a cosine schedule was employed to gradually reduce $\tau_{t}$ to 0.04 over the first 30 epochs. $\lambda_{a}$, $\lambda_{b}$, and $\lambda_{c}$ were assigned values of 0.2, 0.3, and 0.2, respectively. $T_{warmup}$ was set to 50 epochs. The coefficients $\alpha$ and $\beta$ were set to 0.2 and 0.1, respectively. The All experiments were conducted using an NVIDIA GeForce RTX 3090 GPU.
Method All Old New
k-means [27] 15.7 16.6 13.9
GCD [35][CVPR’22] 22.0 26.4 20.1
GPC [48][ICCV’23] 25.2 28.1 22.3
InfoSieve [30][NeurIPS’23] 40.1 50.5 24.2
SimGCD [39][ICCV’23] 32.6 39.0 24.8
5.2 Comparison with the State-of-the-Art methods
Comparision with GCD Methods. The effectiveness of DIG-FACE debiasing is most directly reflected in its ability to maintain high accuracy for old facial expression categories while also achieving excellent results in discovering new facial expression categories. In Tab. 2, we present the comparison with the SOTA GCD method on the basic classes of the RAF-DB dataset  [21]. In terms of old ACC, we achieved an average improvement of 16.2% compared to SimGCD [39]. And the ability to discover new expressions has increased steadily. Moreover, Tab. 3 shows the results on the dataset FerPlus [2]. Although InfoSeive [30] shows marginal improvement in the old class in a certain division (i.e., Surprise, Fear, Disgust, and Anger), our method still achieves significant accuracy gains across all classes, including the new class, demonstrating its effective elimination of implicit bias in learning. Furthermore, Tab. 3 shows performance comparison on the large-scale dataset AffectNet [28]. On this challenging dataset, we both achieved the best new category, old category accuracy. Notably, Due to the excessive number of samples in AffectNet, CMS-GCD [5] struggled to produce effective results. In Tab. 4, we report the comparison on the challenging RAF-DB-Compound dataset that discovers composite expressions based on basic expressions, we achieved an improvement of 20.8% accuracy for the old categories and 4.9% for the new categories, compared to SimGCD. In summary, our approach effectively eliminates the bias towards new categories while maintaining high accuracy for old categories, and steadily improves the recognition accuracy for new categories.
Comparision with FER Method. To verify the generalizability of our method over previous FER approaches, we selected Ada-CAM [19], a semi-supervised method adapted for the G-FACE task, as a baseline. Ada-CAM achieves an average All ACC of only 10.5%. By integrating our DIG-FACE framework into Ada-CAM, we achieve notable improvements of 5.7%, 4.8%, and 2.1% in All ACC across different settings, demonstrating the superior adaptability of our approach. Detailed results are provided in the supplemental appendix. Additionally, adapting to the recent Open-set FER scenario introduced by Zhang et al. [45] et al., we used AUROC (higher is better) and FPR@TPR95 (lower is better) as metrics. The figure is shown in the appendix, DIG-FACE achieved higher AUROC, indicating strong performance in detecting new categories, and lower FPR@TPR95, reflecting greater decision confidence.
Table 5: Ablation studies of implicit and explicit debiasing on RAF-DB, with the four known classes shown above.
Components Sur.+Fea.+Dis.+Hap. Sur.+Fea.+Dis.+Ang.
Implicit Explicit All Old New All Old New
43.3 47.9 40.2 49.6 56.9 48.5
$\checkmark$ 50.4 63.2 40.5 50.5 60.5 48.6
$\checkmark$ 49.8 61.1 40.7 51.5 63.2 50.3
$\checkmark$ $\checkmark$ 53.3 71.3 41.0 52.4 66.3 50.2
Figure 4: Attention visualization of different self-attention heads (numbered as h1 to h3) on RAF-DB. The top 10% attended patches are shown in red. Our method pays more attention to the cheeks, eyes and mouth corners details.
5.3 Ablation Studies
To verify the effectiveness of two debiasing components in DIG-FACE, we conducted ablation experiments on the RAF-DB dataset. The results are presented in Tab. 5 and the more details table in the appendix.
Implicit Debiasing is designed to ensure that learning for known categories remains unaffected by the introduction of new categories. Adding this component to the baseline SimGCD in Tab. 5 resulted in absolute improvements of 15.3% and 3.6% for known categories and relative improvements of 0.75% and 0.21% for new categories. This component focuses on minimizing the bias upper bound, ensuring robust feature extraction even with sub-optimal decision boundaries, thereby bridging the gap between old and new category learning.
Explicit Debiasing. Our explicit debiasing component improves feature separability between known and new categories through Sample-Level and Distribution-Level optimizations, with Triplet-Level contrastive learning adopted from prior GCD work. Incorporating the entire explicit debiasing module leads to absolute improvements of 13.2% and 6.3% for known categories and relative improvements of 1.2% and 3.7% for new categories. Notably, combining the sample-level optimization with implicit debiasing achieves All ACC improvements of 9.0% and 1.1%, while combining the distribution-level optimization with implicit debiasing results in All ACC improvements of 9.8% and 1.3%. Further ablation studies are provided in the appendix.
In general, the combination of implicit and explicit debiasing components leads to the most significant performance improvements. Together, they achieve absolute gains of 23.4% and 9.4% for known categories and relative improvements of 2.0% and 3.5% for new categories, as shown in Tab.5. This demonstrates that implicit debiasing stabilizes learning for known categories, while explicit debiasing enhances feature separability, resulting in complementary benefits for overall model performance.
Figure 5: Compared to the baseline SimGCD, the features extracted by our method are more separable in t-SNE visualization.
Herbarium 19 CIFAR10
Methods All Old New All Old New
$k$-means [27] 13.0 12.2 13.4 83.6 85.7 82.5
UNO+ [10] 28.3 53.7 14.7 68.6 98.3 53.8
GCD [35] 35.4 51.0 27.0 91.5 97.9 88.2
CMS-GCD [5] 36.4 54.9 26.4 94.5 96.0 93.2
SimGCD [39] 44.0 58.0 36.4 97.1 95.1 98.1
DIG-FACE 45.2 58.7 37.9 97.0 95.3 97.8
6 Analysis
6.1 Visualization
As shown in Fig. 4, we adopt the method from GCD [35] to analyze the attention maps generated by the DINO-ViT model. Specifically, we visualize Heads 1 through 3, the self-attention heads in ViT, which reveal how the model captures subtle facial expression details. Compared to GCD and SimGCD, DIG-FACE places greater emphasis on key regions like the cheeks and mouth corners, providing evidence for the improved performance of DIG-FACE. Besides, the t-SNE visualization results shown in Fig. 5 demonstrate a more discrete feature representation learned by DIG-FACE, compared to the more blurred feature boundaries via SimGCD. It shows that DIG-FACE can better reduce the overlap of feature spaces between known and unknown expression classes.
6.2 Performance on General Datasets
In addition to GCD-based FER datasets, we evaluate DIG-FACE on two general image classification datasets, namely CIFAR-10 [17] and Herbarium 19 [32], to assess its generalization ability. Compared to prior GCD methods and the SimGCD baseline, DIG-FACE demonstrates superior robustness and adaptability, as shown in Tab .6.
7 Conclusion
In this paper, we introduced G-FACE, a novel task addressing the dual challenge of recognizing known facial expressions and discovering unseen ones. We defined two key challenges in this task: implicit bias, arising from distributional gaps between labeled and unlabeled data, and explicit bias, caused by shifts in visual facial characteristics between known and unknown categories. To address these challenges, we proposed DIG-FACE, a de-biasing framework that includes implicit debiasing, which minimizes bias through adversarial learning, and explicit debiasing, which enhances feature separability to address category ambiguity. Extensive experiments demonstrate its effectiveness, significantly enhancing recognition for both known and new categories, setting a strong benchmark for G-FACE.
Supplementary Material
8 Experiments
8.1 Implementation Details
We adopt DINO ViT-B/16 as the feature extractor, fine-tuning only the final Transformer blocks, which outputs 768-dimensional representations. The main head maps normalized bottleneck features ( 256-dimensional) to the classification space using a weight-normalized linear layer. The auxiliary head, designed for adversarial training, includes a gradient reversal layer followed by two fully connected layers with 2048 hidden units, GeLU activation, and dropout, enhancing robustness against biases.
8.2 Dataset Details
Table 7: Specific category divisions.
Old Classes New Classes
RAF-DB-Basic $|\mathcal{Y}_{l}|=4$ $|\mathcal{Y}_{u}|-|\mathcal{Y}_{l}|=3$
FerPlus $|\mathcal{Y}_{l}|=4$ $|\mathcal{Y}_{u}|-|\mathcal{Y}_{l}|=4$
AffectNet $|\mathcal{Y}_{l}|=4$ $|\mathcal{Y}_{u}|-|\mathcal{Y}_{l}|=4$
For GCD, we treat four out of the seven basic expressions as known categories, with three different selection methods. Notably, in the stage of discovering compound expressions, we use the basic expressions as known categories. The RAF-DB-Basic dataset is focused on basic expressions, while the RAF-DB-Compound dataset includes more complex expressions. We selected a subset of the basic expressions as old classes for the purpose of our experiments. FERPlus [2] is an extension of the FER2013 dataset [11]. We selected a subset of the categories as old classes to discover the remaining new categories in this dataset. AffectNet [28] is a large-scale facial expression recognition dataset. Similar to the selection strategy for FERPlus, we selected four categories as old classes in AffectNet to discover the remaining expressions.
8.3 A More Detailed Ablation Study
In the main text, we have already conducted ablation experiments on the implicit and explicit components. In this section, we will perform a more detailed ablation of these components, as shown in Tab. 8. When combining sample-level optimization with implicit debiasing, we achieve All ACC improvements of 9.0% and 1.1%, while combining distribution-level optimization with implicit debiasing results in All ACC improvements of 9.8% and 1.3%.
Table 8: Ablation studies of different components of our method on the RAF-DB basic expression classes.
Implicit Debiasing: $\mathcal{L}_{\text{ad}}$, Explicit Debiasing: $\mathcal{L}_{\text{bal}}$, $\mathcal{L}_{\text{cluster}}$
$\mathcal{L}_{\text{ad}}$ $\mathcal{L}_{\text{bal}}$ $\mathcal{L}_{\text{cluster}}$ Sur.+Fea.+Dis.+Hap. Sur.+Fea.+Dis.+Ang.
All Old New All Old New
43.3 47.9 40.2 49.6 56.9 48.5
$\checkmark$ 50.4 63.2 40.5 50.5 60.5 48.6
$\checkmark$ $\checkmark$ 52.3 65.3 40.4 50.7 63.4 47.5
$\checkmark$ $\checkmark$ 53.1 71.4 40.6 50.9 64.1 48.7
$\checkmark$ $\checkmark$ 49.8 61.1 40.7 51.5 63.2 50.3
$\checkmark$ $\checkmark$ $\checkmark$ 53.3 71.3 41.0 52.4 66.3 50.2
Table 9: Ablation study on the impact of using the dynamic adjustment term $e_{i}$ in $\mathcal{L}_{\text{bal}}$ on the RAF-DB dataset.
Sur.+Fea.+Dis.+Hap. Sur.+Fea.+Dis.+Ang.
$\mathcal{L}_{\text{bal}}$ All Old New All Old New
w/o $e_{i}$ 52.3 72.6 38.6 50.8 64.1 48.7
w/ $e_{i}$ 53.3 71.3 41.0 52.4 66.3 50.2
Figure 8: Comparison with open-set FER [45] in the open-set scenario.
Figure 9: Category discovery accuracy (ACC) on RAF-DB. The effect of the hyperparameter $\alpha$.
Table 10: Category discovery accuracy (ACC) on RAF-DB containing only known expressions. We test the FER method Ada-CM under GCD setting. $\triangle$ represents the improvement accuracy.
Sur.+Fea.+Dis.+Hap. Sur.+Fea.+Dis.+Ang. Sur.+Hap.+Sad.+Neu.
Method All Old New All Old New All Old New
@@ 8.3 A More Detailed Ablation Study
Figure 10: Effect of different number of clusters in our DIG-FACE on RAF-DB.
8.4 The Influence of the Hyperparameters
Our method has three main hyperparameters, which are the weight of Loss Functions. We show the results of our model in Fig. 6 on RAF-DB by tuning $\lambda_{a}$, $\lambda_{b}$ and $\lambda_{c}$ with different values. In general our model has good stability for different hyperparameters. We note that if too much weight is applied to the adversarial bias loss, the model performance decreases. Since we only use the adversarial loss to estimate the upper bound of bias, paying excessive attention to this loss can make the model difficult to optimize by other components. In addition $\mathcal{L}_{bal}$ should preferably not be weighted more than the contrastive learning loss $\mathcal{L}_{rep}$, which can lead to limitations in the model’s ability to learn feature representations. Furthermore, $\mathcal{L}_{cluster}$ aims to further improve the expression-discriminability of the feature boundaries and we find that the weight that most improves model recognition is 0.2. Here, we have only roughly obtained an optimal combination of loss weights without conducting a precise analysis. Our main focus was on exploring the effectiveness of the components. The hyperparameters $\alpha$ and $\beta$ are analyzed as shown in Fig. 9 and Fig. 6, taking values of 0.2 and 0.1, respectively. We also investigated the effect of $e_{i}$ in $\mathcal{L}_{bal}$, and Tab. 9 shows results with/without $e_{i}$, illustrating that adaptive changes in smoothness are important.
8.5 Comparision with FER Method
To verify the generalizability of our method over previous FER approaches, we selected Ada-CAM [19], a semi-supervised method adapted for the G-FACE task, as a baseline. Ada-CAM achieves an average All ACC of only 10.5%. By integrating our DIG-FACE framework into Ada-CAM, we achieve notable improvements of 5.7%, 4.8%, and 2.1% in All ACC across different settings, demonstrating the superior adaptability of our approach. Detailed results are provided in the Tab. 10. Additionally, adapting to the recent Open-set FER scenario introduced by [45] et al., we used AUROC (higher is better) and FPR@TPR95 (lower is better) as metrics. Fig. 8 shows the results, DIG-FACE achieved higher AUROC, indicating strong performance in detecting new categories, and lower FPR@TPR95, reflecting greater decision confidence.
8.6 Sensitivity of the number of categories
In this study, we use a parametric classifier approach assuming that the categories are known [39, 37]. In this section, we analyze the effect of different number of categories on ACC. Fig. 10 shows the results, and since FER is a highly imbalanced scenario, increasing the number of output categories of the output header appears as a rise in the old categories. In future work, we will further explore methods that do not have a number of categories a priori.
9 Pseudo Code
We summarize the pipeline of DIG-FACE in Algorithm 1.
Algorithm 1 The proposed DIG-FACE framework
Input: Train set $\mathcal{D}=\{\mathcal{D}_{L}\cup\mathcal{D}_{U}\}$, feature exactor $\psi$, main head $h$, auxiliary head $h_{a}$, train epoch $T$, warm-up epochs $w$.
Output: Trained model parameter $\mathcal{S}$.
Initialize: Load DINO  [4] pre-trained parameters for the backbone.
  for $\text{epoch}=0,\cdots,T-1$ do
     Calculate the predictions of the main head with Eq. (13) and Eq. (9);
     Compute the supervised loss $\mathcal{L}_{\text{rep}}^{s}(\mathcal{D}_{L})$ and self-supervised loss $\mathcal{L}_{\text{rep}}^{u}(\mathcal{D})$;
     Calculate the Adversarial Loss $\mathcal{L}_{ad}$ with Eq. (8);
     if $epoch\geq w$ then
        Feed $\mathcal{D}_{L}$ to $\psi(\cdot)$ and then obtain features;
        Compute the Cluster Loss $L_{cluster}(\psi(\mathcal{D}_{L})$;
     end if
  end for
Algorithm 2 The test stage strategy of DIG-FACE
Input: The test set $\mathcal{D}^{t}=\left\{(\boldsymbol{x}_{j},{y}_{j})\right\}\in\mathcal{X}\times\mathcal{Y}$, trained DIG-FACE model $f_{s}$.
Output: Classification accuracy $\mathrm{ACC}$.
   for $\boldsymbol{x}_{i}\in\mathcal{D}^{t}$ do
      Obtain the feature of $\boldsymbol{x}_{j}$ via $f_{s}(\boldsymbol{x}_{j})$
   end for
10 Theory
10.1 Notation